OpenStudy (anonymous):

PLEASE HELP: Brooks built two similar wooden chests in the shape of rectangular prisms. She used 5 pints of paint to cover the outside of the larger chest fully and 2 pints to cover the outside of the smaller chest. How much more volume does the larger chest contain? A. 158% more B. 250% more C. 300% more D. 395% more

OpenStudy (anonymous):

depends, does the chest have a top?

OpenStudy (anonymous):

another one that is sort of guessing but pretty certain just do 5/2=2.5 then times 100 for 150%

OpenStudy (anonymous):

250% *

OpenStudy (anonymous):

It's not 250%

OpenStudy (anonymous):

2 similar chests meaning the sides of the smaller chest is smaller by a specific ratio which i will denote as "a"

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now, denoting the sides of the chest L x W x H as length width and height now assuming that the chest has a top the surface area of the rectangular prism or chest is SA = 2LW +2LH + 2 WH

OpenStudy (anonymous):

we know the surface area of the larger chest takes 5 pints and the surface area of the smaller chest takes only 2 pints to cover

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

300%

OpenStudy (anonymous):

That's what i got.

OpenStudy (anonymous):

i have no idea, i didnt actually solve it yet

OpenStudy (anonymous):

Haha.

OpenStudy (anonymous):

im kinda solving it as i go

OpenStudy (anonymous):

anyways SA of the smaller chest would be SAsmaller = a^2(2LW +2LH + 2 WH)

OpenStudy (anonymous):

now just assuming that SAlarger = 5 and SAsmaller = 2 substuting them in we get 2=a^2(2LW +2LH + 2 WH) 5=(2LW +2LH + 2 WH)

OpenStudy (anonymous):

now we solve for a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2/5 = a^2

OpenStudy (anonymous):

it wasnt 250 tho. I tried that and got it incorrect.

OpenStudy (anonymous):

i know, its definitely not 250%

OpenStudy (anonymous):

anyways since we know a = sqrt{2/5}

OpenStudy (anonymous):

recall that the volume of a rectangular prism is V= L*W*H

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so the smaller volume would be Vsmall=aL * aW *a H = a^3 LWH and the larger volume would just be Vlarge= L*W*H

OpenStudy (anonymous):

now to find the percentage of increase of area you would need to divide the volume of the large by the volume of the small Vlarge/ V small

OpenStudy (anonymous):

= 1/ a^3

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So it was 158%

OpenStudy (anonymous):

It's either A or C. I think its A.

OpenStudy (anonymous):

tell me, a= ?

OpenStudy (anonymous):

percentage is 158

OpenStudy (anonymous):

i got something different

OpenStudy (anonymous):

rlly? :/

OpenStudy (anonymous):

a = sqrt{2/5}

OpenStudy (anonymous):

a= .63245 ... so the solution would be 1/a^3

OpenStudy (anonymous):

Haha.. okay... so if it isnt A/B/D. Then it must be 300% 282 ?

OpenStudy (anonymous):

why is it not A B D ?

OpenStudy (anonymous):

i got something diffrnt

OpenStudy (anonymous):

Because B i already got wrong. You said 158% was incorrect. and i dont think its 395% because its too high. So what's left is 300%

OpenStudy (anonymous):

dont assume

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

a^3 = .25297557 ...

OpenStudy (anonymous):

so it was 395

OpenStudy (anonymous):

grab a calculator 1/ .25297557 = ?

OpenStudy (anonymous):

which is why you dont assume

OpenStudy (anonymous):

just read through the steps again

OpenStudy (anonymous):

i wasnt assuming. i just didnt round to get the answer. thats why i kept getting 250 as my answer. thanks :)